# PolSAM: Polarimetric Scattering Mechanism Informed Segment Anything Model

### v1

@@ Abstract
Index Terms:
PolSAR terrain segmentation, segment anything model, physical scattering characteristic, prompt-based fusion learning.
@@
I Introduction
Polarimetric Synthetic Aperture Radar (PolSAR) provides rich polarization features, enabling detailed terrain and ground target representation for applications like land-cover segmentation with enhanced accuracy and feature discrimination [1, 2]. Recently, deep learning-based methods for PolSAR image segmentation have gained traction, typically reformatting PolSAR data into specific input types and designing tailored neural networks to extract both image and polarimetric features [3, 4, 5, 6]. Common representations include complex-valued inputs [7], model-based polarimetric features [8, 9, 6], and amplitude images [10], processed using architectures like complex-valued neural networks (CVNNs) [11, 12], convolutional neural networks (CNNs) [3, 6, 4], Transformers [13], and graph neural networks (GNNs) [14, 15]. However, these input formats face challenges in usability, interpretability, and data integrity, while these models are relatively small, limiting their representational capacity.
As illustrated in Fig. 1(a), complex-valued inputs retain the complete information of PolSAR data but require significantly high storage capacity. Furthermore, their processing necessitates the use of specialized CVNNs, which introduces additional computational complexity and weakens interpretability. This constraint also limits the applicability of existing large-scale vision models, as they are typically designed to handle real-valued data, making the integration of complex-valued inputs more challenging. Model-based polarimetric features enhance the interpretability of complex-valued PolSAR data, allowing for effective integration with popular deep learning architectures such as CNNs, Transformers, and GNNs, however, these features are also high-dimensional. The most widely used format is the 8-bit quantized amplitude PolSAR image, aligned with photographic conventions for compatibility with existing deep learning models. Nevertheless, this approach sacrifices the rich polarimetric features, leading to significant information loss. Investigating a user-friendly input format for PolSAR data that retains integrated polarimetric information is vital in deep learning applications.
In PolSAR image segmentation, the challenges discussed above are further intensified by the limited availability of data. This limitation is particularly pronounced in networks with fewer parameters, as they often fail to extract sufficient information from the input data. A common solution is to transfer knowledge from pre-trained models [16], but the high dimensionality of polarimetric features often makes them incompatible with standard image representations [17]. Although CVNNs have twice the number of parameters compared to traditional real-valued networks, the lack of effective pre-trained models limits their utility. In data-scarce scenarios, CVNNs often struggle to converge, further undermining their practical applicability [18, 19]. While various advanced architectures have been proposed to mitigate these challenges, the inherent limitations of small network sizes hinder their ability to effectively extract relevant features. Additionally, the diversity of PolSAR datasets, variations in scattering mechanisms, and differences in imaging conditions complicate the development of models capable of delivering robust performance across diverse scenarios [6].
@@ I Introduction
To address these challenges, this work focuses on designing an efficient input format for PolSAR data that integrates Pauli pseudo-colored images with inherent scattering properties of PolSAR, preserving the core polarimetric features crucial for analysis. Previous studies [13, 20] have demonstrated the critical role of scattering mechanisms in PolSAR analysis. As shown in Fig. 1(b), we optimize the scattering properties into a compact image format, reducing storage requirements while providing intuitive physical interpretability. Moreover, specific scattering patterns basically correspond to particular land cover categories, highlighting a strong correlation with semantic label information. With this foundation, we aim to create a generalized segmentation framework for PolSAR image that leverages the robust capabilities of existing large vision models. This approach seeks to bridge the gap between the unique characteristics of PolSAR data and the good generalization ability of advanced large vision models, enabling broader applications and improved usability for non-expert users.
@@
Building on these insights, we introduce the Segment Anything Model (SAM) [21], as shown in Fig. 1(c). First, SAM leverages its large-scale architecture and robust data processing capabilities to achieve efficient and precise segmentation across diverse image content, making it a promising pre-trained foundation model for PolSAR applications. Second, SAM’s prompt-based segmentation mechanism effectively incorporates high-level external prompts, guiding the model to focus on relevant features. This capability is particularly well-suited for compact, high-level scattering mechanism data, providing semantic guidance to enhance segmentation accuracy.
@@
    We propose a novel processing pipeline for PolSAR data, focusing on generating user-friendly, physically interpretable, and information-dense MVD products. These lightweight data products effectively leverage the strong correlation between scattering mechanisms and semantics to assist in terrain segmentation of PolSAR data.
@@
    Experiments on two newly constructed PhySAR-Seg datasets confirm that PolSAM outperforms existing SAM-based models and multimodal fusion methods. The sparse and dense prompts generated by our approach capture essential semantic relevance, delivering precise guidance for segmentation tasks. Additionally, the integration of MVD reduces data storage requirements, offering a more efficient alternative to conventional polarimetric representations.
II Related Works
II-A Deep learning based PolSAR image analysis
@@
Deep learning techniques have become essential in PolSAR image analysis, leveraging their advanced information processing capabilities to improve segmentation performance [3, 4, 5, 6]. Currently, widely used methods include CNNs or Transformer-based architectures, graph neural networks (GNNs), and complex-valued neural networks (CVNNs). The inputs processed by these methods can be broadly classified into three categories: complex-valued data, model-based polarimetric features, and simple amplitude data.
@@
As shown in Fig. 1(a), ”” denotes an affirmative indicator, ”” denotes a negative indicator. Complex-valued data is mainly processed using CVNNs [7, 11, 12]. Clearly, this approach satisfies information integrity (Inf. int.), but it requires high storage and computational resources, making it not lightweight (Lw), while also lacking physical interpretability (Phys. i.). Polarimetric features (Pol. feature), which mainly include Pauli decomposition-based features [8, 22], polarization coherence matrix $T$ [9, 23, 24, 17], polarization covariance matrix $C$ [6, 25], and scattering features extracted from traditional polarization decomposition methods such as Cloude-Pottier, H/A/Alpha decomposition, Freeman-Durden three-component decomposition, and Yamaguchi four-component decomposition, provide physically interpretable and relatively rich information. These features enhance the ability of deep learning models to interpret PolSAR data [26, 27, 28] . However, they are remain high-dimensional, resulting in high storage and computational costs, making them not lightweight. In comparison, amplitude data, typically represented as 8-bit quantized grayscale images or pseudo-colored images visualized using different polarization modes (HH, HV, VH), is lightweight but suffers from significant information loss and lacks physical interpretability [10]. In summary, these methods face challenges in simultaneously achieving lightweight representation, physical interpretability, and information richness.
@@
Polarimetric features and amplitude data formats can both be processed using networks such as CNNs, Transformers, and GNNs. CNNs and Transformers aim to capture fundamental structural and texture details by enhancing polarization feature representation through advanced techniques [3, 6, 4, 29, 13, 24]. GNNs optimize spatial relationships in PolSAR images by modeling complex spatial dependencies and integrating multi-scale information [30, 27, 31]. Although CNN-based methods effectively utilize polarization features and GNNs capture spatial correlations, these networks often require task-specific designs and training from scratch, limiting their scalability and generalization compared to more general models. Additionally, their computational efficiency still needs improvement.
@@
II-B SAM Implementation in Segmentation
The SAM model has demonstrated considerable success in segmentation tasks, primarily due to its large model size, which enables powerful feature extraction and broad applicability. Additionally, its flexible prompt-based segmentation mechanism allows it to effectively handle external high-level prompt information [32, 33, 34, 35]. This versatility has driven adoption across diverse domains, including natural images [33, 32, 36], medical imaging [34, 35, 37, 38, 39], and remote sensing [40, 41, 42, 43]. In medical imaging and remote sensing, where modality differences necessitate tailored approaches, researchers have adapted SAM with specialized adapter techniques [34, 41, 10]. Further refinements, such as integrating text prompts within segmentation modules, have also enhanced its performance [37]. For natural images, innovations explore the use of a multi-modal encoder that combines visual and linguistic information to generate prompt embeddings [44], while PA-SAM [45] introduces a prompt-driven adapter to enhance SAM’s mask quality by refining sparse and dense prompt features. Altogether, these adaptations highlight SAM’s capacity for domain-specific adjustments, extending its applications across diverse image types.
@@ II-B SAM Implementation in Segmentation
III PhySAR-Seg Dataset
Current PolSAR datasets either rely on quantized amplitude images and pseudo-colored polarization feature images, which lack rich scattering characteristics and physical interpretability, or store complex information, leading to high storage demands. To address these challenges, we propose a novel PolSAR image processing pipeline for constructing the PhySAR-Seg dataset. This dataset is designed to preserve the physical scattering mechanisms in PolSAR images, ensuring both physical interpretability and information integrity, while providing user-friendly, storage-efficient data with strong correlation to meaningful semantic information.
The main innovation in this pipeline is the construction of Microwave Vision Data (MVD), which is defined as a novel representation of PolSAR data. It reflects physical interpretability scattering mechanism and its visual attributes for intuitive understanding. MVD offers several advantages over traditional PolSAR image datasets. First, MVD offers lightweight storage, ensuring high efficiency for large-scale data processing, making it user-friendly. Second, it is both visually and physically interpretable, offering an intuitive visualization of electromagnetic features and a representation of scattering mechanisms with distinct physical meaning. By combining MVD with Pauli pseudo-colored images, we create a more comprehensive dataset. These attributes enhance the efficiency of the PhySAR-Seg dataset, offering deeper insights into both object and terrain characteristics. Details of the processing pipeline and dataset are provided below.
@@ III PhySAR-Seg Dataset
Figure 4: Semantic label clustering of data PhySAR-Seg-1.
@@
III-A PhySAR-Seg Dataset Processing Pipeline
In the first step, the pseudo-colored image is generated from the original single look complex (SLC) data through the Pauli decomposition algorithm, where three components are extracted and assigned to the RGB channels, respectively [46]. Then, the SLC data is processed using the GD-Wishart classification algorithm [47], yielding initial clustering results. The targets are categorized into three classes: odd-bounce, double-bounce, and volume. Notably, in practical experiments, double-bounce scattering is relatively rare, represented by only one class, while the other categories are further divided into five sub-classes each.
To make MVD more closely aligned with semantic information, we re-cluster the scattering classes. Without altering the primary scattering mechanism categories, we re-group the sub-classes within each category based on their semantic correlations. Apart from the three fundamental scattering types, the dataset may include mixed classes and an ’others’ category. The ’others’ category often contains pixels that lack fixed scattering characteristics, typically appearing at boundaries. We then color the classes based on the final clustering results, visualize the data, and save it. This process allows us to generate MVD, which offers a compact yet informative representation of physical scattering mechanisms. MVD efficiently supports large-scale data processing while providing a clear and intuitive visualization of the electromagnetic properties of objects and terrain. By preserving the physical meaning of these scattering mechanisms, MVD not only optimizes storage and processing but also enhances the clear physical interpretability of the dataset for further analysis.
III-B Datasets Description
@@
Based on the Dataset Processing Pipeline outlined above, we constructed two PhySAR-Seg datasets, each comprising pairs of pseudo-colored and MVD images. These datasets are lightweight, physical interpretable, with scattering mechanisms that have clear physical meanings and exhibit certain correlations with semantic information. Their comprehensive representation capabilities provide a solid foundation for evaluating our proposed method, ensuring that the segmentation model can be effectively trained and assessed on diverse and meaningful data.
@@
The PhySAR-Seg-1 dataset is sourced from [48]. For our experiments, we use the L, S, C, and X bands, all covering the same geographic region. The L, S, and C bands have a resolution of 0.5 m, while the X band has a resolution of 1.0 m. Fig. 3 provides an overview of the dataset, showcasing the pseudo-colored images, MVD, and label maps, along with both the MVD and semantic legends. In the MVD legend, colors within the same hue represent consistent scattering types, with lighter shades corresponding to higher SPAN values. The semantic legend displays the proportions of the five final terrain classes, clearly illustrating their distribution across the dataset. Initially, the dataset contains 14 land cover types, which we merge into five categories: urban surfaces, water, buildings, barren land, and vegetation, as shown in Fig. 4. A brief note is that we calculate the MVD for each of the four bands. Upon comparison, the X band shows the best performance. Since all bands cover the same region and capture similar scattering mechanisms, we register the X-band MVD to the other three bands to generate the final result. The segmented dataset comprises 2,866 image pairs of pseudo-colored MVD, each with 512 $\times$ 512 pixels. We divide the data from each band into training, validation, and test sets in a 6:2:2 ratio, following the dashed lines shown in Fig. 3. This approach ensures that data from the same geographic region in different bands is allocated to the same set, preventing any overlap between the training, validation, and test sets.
@@
The PhySAR-Seg-2 dataset is sourced from [49] and is based on spaceborne C-band data. We also obtain MVD according to the method in Fig. 2, and labels are obtained through manual annotation, resulting in a total of seven categories excluding the background. Fig. 5 provides an overview of the dataset, including pseudo-colored images, MVD, and their corresponding semantic labels. Due to the limited number of samples in the dataset, we only partition the data into a training set and a validation set in a 6:4 ratio, as indicated by the dashed line in the figure. Given the substantial presence of water in the dataset, we strategically remove pure water samples from the segmented 512 $\times$ 512 images to maintain a balance among categories. After this adjustment, 1,110 images remain, with 668 allocated to the training set and 442 to the validation set. The semantic legend clearly indicates the category distribution before and after processing.
@@
IV The Proposed Method
IV-A Revisiting SAM
The SAM introduces an innovative prompt-based architecture for diverse segmentation tasks. It consists of three main components: the image encoder ($f_{\text{Enc}}$), the prompt encoder ($f_{\text{PE}}$), and the mask decoder ($f_{\text{Dec}}$). The image encoder extracts high-dimensional feature maps from the input image, which serve as the foundation of SAM’s segmentation capability. The prompt encoder plays a key role by accepting flexible inputs, such as sparse prompts ($I_{\text{sparse}}$, e.g., points or bounding boxes), dense prompts ($I_{\text{dense}}$, e.g., masks), and text. These prompts guide the model to focus on specific regions of interest and are transformed into embeddings that integrate with the image features. Finally, the mask decoder fuses both the sparse and dense embeddings with the image features to produce the final segmentation mask. This approach enables SAM to handle diverse segmentation challenges effectively without task-specific fine-tuning. It is a versatile and robust solution for a wide range of scenarios. This can be mathematically represented as follows:

### v2

@@ Abstract
keywords:
PolSAR terrain segmentation, prompt-based fusion learning, segment anything model, physical scattering characteristic.
1 Introduction
Polarimetric Synthetic Aperture Radar (PolSAR) provides rich polarization features that enable detailed representation of terrain and ground targets, offering enhanced accuracy and feature discrimination for applications such as land-cover segmentation Tirandaz et al. (2020); Bi et al. (2020); Wang et al. (2022). Recently, deep learning-based methods for PolSAR image segmentation have gained significant attention. These approaches typically reformat PolSAR data into specific input representations and design specialized neural networks to extract both spatial and polarimetric features Liu et al. (2018); Zhang et al. (2020b); Qin et al. (2020); Ghanbari et al. (2023). Common input representations include complex-valued data Li et al. (2020), model-based polarimetric features Antropov et al. (2013); Shen et al. (2020); Ghanbari et al. (2023), and amplitude images Pu et al. (2024). These are processed using various architectures, including complex-valued neural networks (CVNNs) Yu et al. (2022); Kuang et al. (2024), convolutional neural networks (CNNs) Liu et al. (2018); Ghanbari et al. (2023); Zhang et al. (2020b), Transformers Hua et al. (2024), and graph neural networks (GNNs) Dong et al. (2021); Jamali et al. (2023). However, these input formats present challenges in terms of usability, interpretability, and data integrity. Moreover, the existing models are typically of limited scale, which constrains their representational capacity.
As illustrated in Fig. 1(a), complex-valued inputs preserve the complete information of PolSAR data but require substantially high storage capacity. Furthermore, their processing necessitates the use of specialized CVNNs, which introduces additional computational complexity and reduces interpretability. This constraint also limits the applicability of existing large-scale vision models, as they are primarily designed to handle real-valued data, making the integration of complex-valued inputs challenging. Model-based polarimetric features enhance the interpretability of complex-valued PolSAR data, enabling effective integration with popular deep learning architectures such as CNNs, Transformers, and GNNs; however, these features are also high-dimensional. The most widely used format is the 8-bit quantized amplitude PolSAR image, which aligns with photographic conventions for compatibility with existing deep learning models Zhang et al. (2020a); Liu et al. (2022); Zhang et al. (2023b). Nevertheless, this approach sacrifices rich polarimetric features, resulting in significant information loss. Investigating a user-friendly input format for PolSAR data that retains comprehensive polarimetric information is essential for deep learning applications.
In PolSAR image segmentation, these challenges become more severe due to limited data availability. This constraint particularly affects networks with fewer parameters, which often fail to extract sufficient information from input data. Knowledge transfer from pre-trained models offers a common solution Wu et al. (2019), yet the high dimensionality of polarimetric features frequently renders them incompatible with standard image representations Huang et al. (2018). Although CVNNs possess twice the parameters of traditional real-valued networks, the absence of effective pre-trained models constrains their utility. In data-scarce scenarios, CVNNs frequently encounter convergence difficulties, further limiting their practical applicability Zeng et al. (2022, 2023). Despite various advanced architectures proposed to address these challenges, the inherent constraints of small network architectures impede effective feature extraction. Moreover, the diversity of PolSAR datasets, variations in scattering mechanisms, and differences in imaging conditions complicate the development of models capable of robust performance across diverse scenarios Ghanbari et al. (2023).
@@ 1 Introduction
To address these challenges, this work designs an efficient input format for PolSAR data that integrates Pauli pseudo-colored images with inherent scattering properties, preserving essential polarimetric features for analysis. Previous studies Hua et al. (2024); Huang et al. (2022) have established the critical importance of scattering mechanisms in PolSAR analysis. As illustrated in Fig. 1(b), we optimize scattering properties into a compact image format that reduces storage requirements while maintaining intuitive physical interpretability. Notably, specific scattering patterns correspond directly to particular land cover categories, revealing strong correlations with semantic label information. Building on this foundation, we develop a generalized segmentation framework for PolSAR images that harnesses the robust capabilities of existing large vision models. This approach bridges the gap between PolSAR data’s unique characteristics and the superior generalization ability of advanced large vision models, enabling broader applications and enhanced usability for non-expert users.
@@
Building on these insights, we introduce the Segment Anything Model (SAM) Kirillov et al. (2023), as shown in Fig. 1(c). First, SAM leverages its large-scale architecture and robust data processing capabilities to achieve efficient and precise segmentation across diverse image content, making it a promising pre-trained foundation model for PolSAR applications. Second, SAM’s prompt-based segmentation mechanism effectively incorporates high-level external prompts, guiding the model to focus on relevant features. This capability is particularly well-suited for compact, high-level scattering mechanism data, providing semantic guidance to enhance segmentation accuracy.
@@
    We propose a novel pipeline for PolSAR data that generates user-friendly, physically interpretable, and information-dense MVD products. These lightweight products leverage the strong correlation between scattering mechanisms and semantics to aid in terrain segmentation.
@@
    Experiments on the PhySAR-Seg datasets show that PolSAM outperforms SAM-based and multimodal fusion models. The sparse and dense prompts effectively capture semantic relevance, guiding segmentation tasks. Additionally, MVD integration reduces storage requirements, offering a more efficient alternative to traditional polarimetric representations.
2 Related Works
2.1 Deep learning based PolSAR image analysis
@@
Deep learning techniques have become essential in PolSAR image analysis, leveraging their advanced information processing capabilities to improve segmentation performance Liu et al. (2018); Zhang et al. (2020b); Qin et al. (2020); Ghanbari et al. (2023). Currently, widely used methods include CNNs or Transformer-based architectures, GNNs, and CVNNs. The inputs processed by these methods can be broadly classified into three categories: complex-valued data, model-based polarimetric features, and simple amplitude data.
@@
As shown in Fig. 1(a), "" denotes an affirmative indicator, "" denotes a negative indicator. Complex-valued data is mainly processed using CVNNs Li et al. (2020); Yu et al. (2022); Kuang et al. (2024). Clearly, this approach satisfies information integrity (Inf. int.), but it requires high storage and computational resources, making it not lightweight (Lw), while also lacking physical interpretability (Phys. i.). Polarimetric features (Pol. feature), which mainly include Pauli decomposition-based features Antropov et al. (2013), polarization coherence matrix $T$ Shen et al. (2020); Gao et al. (2023); Zhang et al. (2023d), polarization covariance matrix $C$ Ghanbari et al. (2023); Xiang et al. (2024), and scattering features extracted from traditional polarization decomposition methods such as Cloude-Pottier, H/A/Alpha decomposition, Freeman-Durden three-component decomposition, and Yamaguchi four-component decomposition, provide physically interpretable and relatively rich information. These features enhance the ability of deep learning models to interpret PolSAR data Xiao et al. (2020); Shi et al. (2023) . However, they are remain high-dimensional, resulting in high storage and computational costs, making them not lightweight. In comparison, amplitude data, typically represented as 8-bit quantized grayscale images or pseudo-colored images visualized using different polarization modes (HH, HV, VH), is lightweight but suffers from significant information loss and lacks physical interpretability Pu et al. (2024). In summary, these methods face challenges in simultaneously achieving lightweight representation, physical interpretability, and information richness.
@@
Polarimetric features and amplitude data formats can both be processed using networks such as CNNs, Transformers, and GNNs. CNNs and Transformers aim to capture fundamental structural and texture details by enhancing polarization feature representation through advanced techniques Liu et al. (2018); Ghanbari et al. (2023); Zhang et al. (2020b); Hua et al. (2024); Zhang et al. (2023d). GNNs optimize spatial relationships in PolSAR images by modeling complex spatial dependencies and integrating multi-scale information Geng et al. (2022); Shi et al. (2023); Wang et al. (2024). Despite the aforementioned advancements, these networks often require task-specific designs and their limited scale restricts scalability and computational efficiency.
@@
2.2 SAM Implementation in Segmentation
The SAM model has demonstrated considerable success in segmentation tasks, primarily due to its large model size, which enables powerful feature extraction and broad applicability. Additionally, its flexible prompt-based segmentation mechanism allows it to effectively handle external high-level prompt information Zhang et al. (2023e, a); Wu et al. (2023); Chai et al. (2023). This versatility has driven adoption across diverse domains, including natural images Zhang et al. (2023a, e); Ke et al. (2024), medical imaging Wu et al. (2023); Chai et al. (2023); Yue et al. (2023, 2024); Xu et al. (2024b, a), and remote sensing Chen et al. (2024); Zhang et al. (2024a); Yan et al. (2023); Zheng et al. (2024); Ma et al. (2024); Zhou et al. (2024). In medical imaging and remote sensing, where modality differences necessitate tailored approaches, researchers have adapted SAM with specialized adapter techniques Wu et al. (2023); Zhang et al. (2024a); Pu et al. (2024). Further refinements, such as integrating text prompts within segmentation modules, have also enhanced its performance Yue et al. (2023). For natural images, innovations explore the use of a multi-modal encoder that combines visual and linguistic information to generate prompt embeddings Zhang et al. (2024b), while PA-SAM Xie et al. (2024) introduces a prompt-driven adapter to enhance SAM’s mask quality by refining sparse and dense prompt features. Altogether, these adaptations highlight SAM’s capacity for domain-specific adjustments, extending its applications across diverse image types.
@@ 2.2 SAM Implementation in Segmentation
3 PhySAR-Seg Dataset
To tackle the data challenges mentioned above, we propose a novel pipeline for constructing the PhySAR-Seg dataset. The key innovation is Microwave Vision Data (MVD), a lightweight and user-friendly representation that ensures high storage efficiency and provides intuitive, physically interpretable visualizations of scattering mechanisms. By integrating MVD with Pauli pseudo-colored images, we also ensure information integrity. Details of the pipeline and dataset are provided below.
3.1 PhySAR-Seg Dataset Processing Pipeline
In the first step, we generate pseudo-colored images from original single look complex (SLC) data using the Pauli decomposition algorithm, which extracts three components and assigns them to RGB channels respectively Huang et al. (2021). Subsequently, the SLC data undergoes processing with the GD-Wishart classification algorithm Ratha et al. (2017) to produce initial clustering results. Targets are categorized into three primary classes: odd-bounce, double-bounce, and volume scattering. In practical experiments, double-bounce scattering occurs infrequently and is represented by a single class, while odd-bounce and volume categories are each subdivided into five sub-classes.
To better align MVD with semantic information, we re-cluster scattering classes by grouping sub-classes within each primary category based on semantic correlations. Beyond the three primary scattering types, the dataset incorporates mixed and ’other’ classes, with the latter typically containing boundary pixels of undefined scattering characteristics. Through color-coding and cluster visualization, we generate MVD as a compact and informative representation of scattering mechanisms. This representation enables efficient large-scale processing, provides intuitive electromagnetic property visualization, and preserves physical meaning while optimizing storage requirements and enhancing interpretability for subsequent analysis.
Figure 3: Semantic label clustering of data PhySAR-Seg-1.
@@ 3.1 PhySAR-Seg Dataset Processing Pipeline
3.2 Datasets Description
@@
Based on this processing pipeline, we construct two lightweight, physically interpretable, and information-integrated PhySAR-Seg datasets, each comprising paired pseudo-colored and MVD images. Their comprehensive representation capabilities provide a solid foundation for evaluating our proposed method.
@@
The PhySAR-Seg-1 dataset derives from Yan et al. (2024) and includes L, S, C, and X bands covering the same geographic region. The L, S, and C bands achieve 0.5 m resolution, while the X band provides 1.0 m resolution. Fig. 4 presents a dataset overview, displaying pseudo-colored images, MVD, and label maps alongside both MVD and semantic legends. In the MVD legend, colors within identical hues represent consistent scattering types, with lighter shades indicating higher SPAN values. The semantic legend shows proportions of five final terrain classes, clearly illustrating their distribution. Initially containing 14 land cover types, we merge these into five categories: urban surfaces, water, buildings, barren land, and vegetation, as shown in Fig. 3. We calculate MVD for each band, with comparative analysis revealing optimal X-band performance. Since all bands cover identical regions and capture similar scattering mechanisms, we register X-band MVD to the remaining three bands for final result generation. The segmented dataset contains 2,866 pseudo-colored MVD image pairs, each measuring 512 × 512 pixels. We partition data from each band into training, validation, and test sets using a 6:2:2 ratio, following the dashed lines in Fig. 4. This strategy ensures that data from identical geographic regions across different bands remains within the same set, preventing overlap between training, validation, and test partitions.
@@
The PhySAR-Seg-2 dataset derives from Gaofen-3 (2019) and utilizes spaceborne C-band data. We obtain MVD following the methodology in Fig. 2, while labels are acquired through manual annotation, yielding seven categories excluding background. Fig. 5 presents the dataset overview, displaying pseudo-colored images, MVD, and corresponding semantic labels. Due to limited sample availability, we partition the data into training and validation sets using a 6:4 ratio, as indicated by the dashed line in the figure. Given the substantial water presence in the dataset, we strategically remove pure water samples from the segmented 512 × 512 images to maintain categorical balance. Following this adjustment, 1,110 images remain, with 668 assigned to training and 442 to validation. The semantic legend clearly illustrates category distribution before and after processing.
@@
4 The Proposed Method
4.1 Revisiting SAM
SAM introduces an innovative prompt-based architecture for diverse segmentation tasks, comprising three main components: the image encoder ($f_{\text{Enc}}$), prompt encoder ($f_{\text{PE}}$), and mask decoder ($f_{\text{Dec}}$). The image encoder extracts high-dimensional feature maps from input images, establishing the foundation for SAM’s segmentation capabilities. The prompt encoder serves a pivotal role by processing flexible inputs including sparse prompts ($I_{\text{sparse}}$, such as points or bounding boxes), dense prompts ($I_{\text{dense}}$, such as masks), and text. These prompts direct the model toward specific regions of interest and are transformed into embeddings that integrate with image features. The mask decoder subsequently fuses both sparse and dense embeddings with image features to generate final segmentation masks. This architecture enables SAM to address diverse segmentation challenges effectively without requiring task-specific fine-tuning, providing a versatile and robust solution across various scenarios. The process can be mathematically represented as follows: